### v1

# Graph Positional Autoencoders as Self-supervised Learners

Abstract.
Graph self-supervised learning seeks to learn effective graph representations without relying on labeled data. Among various approaches, graph autoencoders (GAEs) have gained significant attention for their efficiency and scalability. Typically, GAEs take incomplete graphs as input and predict missing elements, such as masked nodes or edges. While effective, our experimental investigation reveals that traditional node or edge masking paradigms primarily capture low-frequency signals in the graph and fail to learn the expressive structural information. To address these issues, we propose Graph Positional Autoencoders (GraphPAE), which employs a dual-path architecture to reconstruct both node features and positions. Specifically, the feature path uses positional encoding to enhance the message-passing processing, improving GAE’s ability to predict the corrupted information. The position path, on the other hand, leverages node representations to refine positions and approximate eigenvectors, thereby enabling the encoder to learn diverse frequency information. We conduct extensive experiments to verify the effectiveness of GraphPAE, including heterophilic node classification, graph property prediction, and transfer learning. The results demonstrate that GraphPAE achieves state-of-the-art performance and consistently outperforms baselines by a large margin.
Graph Neural Networks, Self-supervised Learning, Graph Autoencoders, Positional Encoding
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3736990ccs: Information systems Data mining
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15516721.
1. Introduction
Table 1. Comparison of different graph autoencoders.
Model Corruption Reconstruction
Feature Edge Position Feature Edge Other
GraphMAE (Hou et al., 2022) - - - -
StructMAE (Liu et al., 2024b) - - - -
AUG-MAE (Wang et al., 2024) - - - -
S2GAE (Tan et al., 2023) - - - -
SeeGera (Li et al., 2023b) - -
Bandana (Zhao et al., 2024) - - - -
MaskGAE (Li et al., 2023a) - - - Degree
GiGaMAE (Shi et al., 2023) - - - Latent
GraphPAE - - Position
(a) Feature Masking
(b) Edge Masking
(c) Eigenvector offset
Figure 1. Frequency magnitude comparison between the original graph and its corrupted versions.
Graph neural networks (GNNs) have achieved significant success across various fields, including social network analysis (Carrington et al., 2005; Singh et al., 2024; Yu et al., 2025), recommendation systems (Gao et al., 2023; Sharma et al., 2024), and drug discovery (Jiang et al., 2021; Bongini et al., 2021; Xiong et al., 2019). However, training effective GNNs in real-world applications remains challenging due to the limited availability of labeled data in many domains (Yu et al., 2024a; Zhu et al., 2025). To address this problem, graph self-supervised learning is proposed to learn graph representations without labeled data (Liu et al., 2022; Xie et al., 2022; Yu et al., 2024b).
Among existing approaches, the generative (Wu et al., 2021b; Waikhom and Patgiri, 2023) and contrastive (Ju et al., 2024) learning paradigms have dominated recent advances. In particular, graph autoencoders (GAEs) (Kipf and Welling, 2016; Hou et al., 2022) have regained attention due to their simplicity, efficiency, and scalability. The GAEs follow a corruption-reconstruction framework, which learns graph representations by recovering the missing information of the incomplete input graphs, as Table 1 illustrates. For example, GraphMAE (Hou et al., 2022) replaces node features with a learnable token, while Bandana (Zhao et al., 2024) proposes a non-discrete edge masking strategy. Moreover, some GAEs even go beyond reconstructing nodes and edges by targeting structural features like degree (Li et al., 2023a). Despite their success, the performance of GAEs highly depends on the choice of corruption and reconstruction objectives. Therefore, a natural question arises: Do existing feature or edge masking mechanisms enable GAEs to sufficiently capture key structural patterns in graph data? If not, what alternative objectives could be designed to further improve the performance of GAEs? A well-informed answer can not only help us identify the weaknesses of existing masking strategies but also deepen our understanding of GAEs.
To answer this question, we first revisit the feature and edge masking strategies from a spectral perspective (Bo et al., 2023b, a). Specifically, we transform the node features $\mathbf{X}$ into the spectral domain by using the eigenvectors of graph Laplacian $\mathbf{U}^{\top}$ and examine its frequency magnitude $\mathbf{U}^{\top}\mathbf{X}$. In the case of feature masking, we randomly mask 20% node features in the Squirrel dataset, denoting as $\widetilde{\textbf{X}}$. The comparison between $\mathbf{U}^{\top}\mathbf{X}$ and $\mathbf{U}^{\top}\widetilde{\textbf{X}}$ is shown in Figure 1(a). For edge masking, we randomly remove 20% edges and construct the corrupted graph eigenvectors $\widetilde{\textbf{U}}^{\top}$. Figure 1(b) compares $\mathbf{U}^{\top}\mathbf{X}$ and $\widetilde{\textbf{U}}^{\top}\mathbf{X}$. Experimental details are provided in Appendix A.1. From these comparisons, we observe that the magnitude differences between original and corrupted graphs are more pronounced in the low-frequency band [0.0-0.1], while the differences decrease at higher frequencies. This indicates that existing GAEs focus primarily on reconstructing low-frequency information. Specifically, the significant difference in the [0.0-0.1] band suggests that masking predominantly damage these low-frequency components. During reconstruction, the GAE is optimized by minimizing the reconstruction loss, which inherently penalizes larger discrepancies more heavily to accelerate convergence. Since low-frequency components are corrupted more severely, GAE implicitly tends to prioritize these components to reduce the loss quickly. As a result, existing GAEs often overlook high-frequency information, which has been shown to be valuable in real-world tasks, such as heterophilic node classification (Bo et al., 2021; Wang et al., 2020).
Once the weakness of existing GAEs is identified, it is natural to ask: How can we design the corruption and reconstruction objectives to enable GAEs to learn diverse frequency information? Essentially, the eigenvectors of the graph Laplacian represent different frequencies. Motivated by this, a straightforward approach is to perturb eigenvectors. Figure 1(c) illustrates the impact on frequency magnitudes of corrupted eigenvectors $\widetilde{\mathbf{U}}$ with offsets. By comparing $\mathbf{U}^{\top}\mathbf{X}$ and $\widetilde{\mathbf{U}}^{\top}\mathbf{X}$, we observe that the eigenvector offset corrupts higher-frequency signals and covers a broader range. These observations indicate that the eigenvector offset introduces wider perturbations than masking, enabling the model to capture more diverse information. However, this task is non-trivial and presents two main challenges: (1) Eigenvectors represent the underlying structures of a graph, which cannot be easily approximated by basic GNNs. Existing GAEs typically adopt message-passing neural networks (MPNNs) as the encoder, whose expressiveness is bounded by the 1-WL test (Xu et al., 2018). The intrinsic weakness of MPNNs restricts GAEs’ ability to capture long-range dependencies between nodes and higher-order graph patterns (Wu et al., 2021a; Dwivedi et al., 2022). (2) Eigenvectors suffer from sign- and basis-ambiguity issues (Lim et al., 2022). Directly reconstructing the eigenvectors leads to non-unique solutions, affecting the robustness of GAEs (Wang et al., 2022; Bo et al., 2024).
To overcome these challenges, we propose Graph Positional Autoencoders (GraphPAE) for graph self-supervised learning. GraphPAE adopts a dual-path architecture to address both the expressivity and ambiguity issues. Specifically, in the feature path, GraphPAE integrates positional encoding (PE) into the message-passing process to enhance the expressiveness of MPNNs (Dwivedi et al., 2023; Wang et al., 2022), thus improving the model’s ability to reconstruct corrupted information. Moreover, in the position path, node representations are used to refine the PE, which approximates the eigenvectors, thereby transferring diverse frequency information into the encoder. Finally, in the reconstruction stage, instead of directly recovering the raw eigenvectors, GraphPAE uses the relative node distance as a surrogate objective, avoiding potential ambiguity issues.
Contributions. The contributions of our paper are as follows:
  1. (1)
    To the best of our knowledge, we are the first to explore the masking strategy in GAEs from a spectral perspective. By comparing the frequency magnitudes between original and corrupted graphs, we identify that existing GAEs focus on reconstructing the extremely low-frequency information of graphs and neglecting other frequencies.
  2. (2)
    We propose GraphPAE, a novel graph positional autoencoder that reconstructs both node features and positions, thereby enabling GAEs to capture a broader range of frequency information. GraphPAE adopts a dual-path architecture. Specifically, the feature path generates expressive node presentations enhanced by positional encodings, and the position path outputs positional encodings to facilitate position reconstruction.
  3. (3)
    We benchmark GraphPAE against state-of-the-art baselines across various tasks, including node classification, graph property prediction, and transfer learning. The results on 14 graphs demonstrate that GraphPAE consistently outperforms the baselines and shows impressive performance on heterophilic graphs.
2. Related Work
2.1. Generative Graph Self-supervised Learning
Generative graph learning encompasses a set of graph self-supervised learning techniques aimed at reconstructing missing information in incomplete input graphs. It can be broadly categorized into autoregressive and autoencoding approaches.
Graph Autoregressive models (GARs) treat sequential graph generation as the pre-training task, where the node or edge is predicted based on its prior context. GPT-GNN (Hu et al., 2020a) factorizes each node generation into attribute and edge generation and replenishes the omitted parts by an adaptive queue. MGSSL (Zhang et al., 2021a) introduces motif generation into pre-training. GraphsGPT (Gao et al., [n. d.]) transforms non-Euclidean graphs into learnable Euclidean words and reconstructs the original graph from these words.
Graph Autoencoders (GAEs) reconstruct all desired content once from the latent representation output by the encoder. Early GAEs (e.g., VGAE (Kipf and Welling, 2016), ARGVA (Pan et al., 2018)) learn representations through link reconstruction and spark a series of work, including feature reconstruction (e.g., GALA (Park et al., 2019), WGDN (Cheng et al., 2023)) and combination reconstruction of structure and feature (e.g., GATE (Salehi and Davulcu, 2019)). However, these traditional GAEs often perform poorly in downstream tasks, except link prediction, which is attributed to their overemphasis on proximity information at the expense of structural information (Hassani and Khasahmadi, 2020). Recently, masked GAEs (Hu et al., [n. d.]; Hou et al., 2022) have become highly successful models for representation learning. GraphMAE (Hou et al., 2022) and GraphMAE2 (Hou et al., 2023) successfully bridge the performance gap between graph contrastive learning and generative learning by reconstructing masked features for training. AUG-MAE (Wang et al., 2024) introduces an adversarial masking strategy to enhance feature alignment and add a uniformity regularizer to promote high-quality graph representations. Additionally, some works are focusing on masking and reconstructing graph structures (e.g., edges (Li et al., 2023a; Tan et al., 2023) and paths (Li et al., 2023a)). Bandana (Zhao et al., 2024) uses continuous and dispersive edge masks and bandwidth prediction instead of discrete edge masks and reconstruction. Besides, some works mask features and edges simultaneously (Li et al., 2023b; Shi et al., 2023) and propose novel reconstruction objectives such as latent embeddings (Shi et al., 2023).
2.2. Graph Positional Encoding
Positional encodings (PEs), originally designed to enhance transformers by encoding positional information in sequential data, have been introduced in graph learning to provide explicit position-aware features. By assigning unique identifiers to nodes based on their structural positions, PEs enable graph neural networks (GNNs) to distinguish non-isomorphic structures beyond the limitations of the 1-WL test. Existing graph PEs can be divided into two main categories: Laplacian-based PEs and Distance-based PEs.
Laplacian-based PEs leverage the eigenvectors of the graph Laplacian matrix as initial node positional features (Dwivedi and Bresson, 2020; Dwivedi et al., 2023, 2021). Eigenvectors form a natural basis for encoding graph structure, similar to how sinusoidal functions represent positional information in sequential models (Vaswani, 2017). Serving as PEs, eigenvectors have two traditional constraints, i.e., sign- and basis ambiguity (Lim et al., 2022). To keep the models robust to sign ambiguity, (Dwivedi and Bresson, 2020; Dwivedi et al., 2023) randomly flip the sign of the eigenvectors during training. SAN (Kreuzer et al., 2021) introduces the information of eigenvalues into PEs and leverages a transformer-based positional encoder, enabling models to be more informative and expressive. PEG (Wang et al., 2022) solves the sign- and basis-ambiguity by treating the distance of eigenvectors between node pairs as PEs. (Lim et al., 2022) propose SignNet and BasisNet to learn sign- and basis-invariant PEs, respectively. (Huang et al., 2023; Bo et al., 2024) study stable PEs that are robust to disturbance of the Laplacian matrix.
Distance-based PEs assign positional features to nodes based on their relative distances within the graph structure, which are typically derived from spatial relationships (Rampášek et al., 2022; Li et al., 2020). One common approach is to use the random walk matrix capturing structural relationships in a probabilistic manner (Dwivedi et al., 2021, 2023). Another popular method is to encode the shortest distance between node pairs (Ying et al., 2021; Rampášek et al., 2022; Li et al., 2020). For example, Graphormer (Ying et al., 2021) incorporates shortest-path distances into its attention mechanism. Additionally, GraphiT (Mialon et al., 2021) encodes PEs with a diffusion kernel, enabling more flexible and adaptive representations of node relationships. (Shiv and Quirk, 2019) introduces a novel positional encoding scheme that extends transformers to tree-structured data, enabling efficient and parallelizable sequence-to-tree, tree-to-sequence, and tree-to-tree mappings. (Wijesinghe et al., [n. d.]) first introduces PEs in graph self-supervised learning and designs a GNN framework that integrates a k-hop message-passing mechanism to enhance its expressiveness.
(a) Overall pipeline of GraphPAE.
(b) Encoder of GraphPAE
Figure 2. (a): GraphPAE integrates a positional corruption-reconstruction mechanism to encourage the GAE to capture diverse frequency information. For feature reconstruction, masked features are encoded with $\mathbf{P}$ and decoded to recover original features. For positional reconstruction, noise is added to $\widetilde{\mathbf{U}}$ to corrupt relative distances, which are then encoded with $\mathbf{X}$ and decoded to recover the original distances. Figure (b): The encoder employs a dual-path architecture to update node $\mathbf{X}^{\left(l\right)}$ and positional representations $\mathbf{P}^{\left(l\right)}$ at each layer. The feature path integrates positional encodings to enhance message passing, improving GAE’s ability to reconstruct corrupted features. The position path utilizes node representations to refine positional embeddings to approximate original pairwise distances.
3. Preliminaries
Before describing our method, we first define some notations and introduce some important concepts used in this paper.
Problem Definition. Given a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$, where $\mathcal{V}$ is the node set with $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$, $\mathcal{E}$ is the set of edges, and $\mathbf{X}\in\mathbb{R}^{N\times d}$ represents the $d$-dimensional node feature matrix. We use the adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$ to describe the graph structure, where $A_{ij}=1$ if there is en edge between $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise. The goal of our framework is to learn an effective graph encoder $f_{enc}(\cdot)$ without relying on labels from downstream tasks. Once trained, the encoder $f_{enc}(\cdot)$ generates node representations as $\mathbf{H}=f_{enc}(\mathcal{G})\in\mathbb{R}^{N\times d_{h}}$ or graph representations as $\mathbf{H}_{\mathcal{G}}=\text{READOUT}\left(\mathbf{H}_{v_{i}}\mid v_{i}\in\mathcal{V}\right)\in\mathbb{R}^{d_{h}}$, where READOUT is a permutation-invariant function such as mean, max, or sum pooling. The representations can then be used to train a simple linear classifier on labeled data from downstream tasks.
Laplacian Graph Eigenvectors. The normalized graph Laplacian matrix $\mathbf{L}$ is defined as $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$, where $\mathbf{I}_{N}$ is an identity matrix and $\mathbf{D}$ is the degree matrix with $D_{ii}=\sum_{j}A_{ij}$ for $v_{i}\in\mathcal{V}$, and $D_{ij}=0$ for $i\neq j$. The Laplacian matrix $\mathbf{L}$ can be decomposed as $\mathbf{L}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$, where $\bm{\Lambda}=\text{diag}(\{\lambda_{i}\}_{i=1}^{N})$ is the diagonal matrix of eigenvalues, and $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{N}]\in\mathbb{R}^{N\times N}$ consists of a set of eigenvectors. Each eigenvector $\mathbf{u}_{i}\in\mathbb{R}^{N}$ corresponds to eigenvalue $\lambda_{i}$, where the eigenvalues are ordered as $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$. Given eigen-decomposition $\mathbf{L}\mathbf{u}_{i}=\lambda_{i}\mathbf{u}_{i}$ with $\mathbf{u}_{i}^{\top}\mathbf{u}_{i}=1$, we have $\mathbf{u}_{i}^{\top}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$. Since $\mathbf{u}_{i}^{\top}\mathbf{L}\mathbf{u}_{i}=\sum_{(j,k)\in\mathcal{E}}(\mathbf{u}_{i,j}-\mathbf{u}_{i,k})^{2}$, we finally derive that $\sum_{(j,k)\in\mathcal{E}}(\mathbf{u}_{i,j}-\mathbf{u}_{i,k})^{2}=\lambda_{i}$, which indicates that $\lambda_{i}$ reflects the frequency magnitude of eigenvector $\mathbf{u}_{i}$ over the graph.
In this paper, we follow the popular positional encoding methods and adopt top-$k$ eigenvectors as initial positions to reduce complexity. Therefore, we redefine $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{K}]\in\mathbb{R}^{N\times K}$, where $K\leq N$. Notably, these eigenvectors are not unique, as sign flips or coordinate rotations may still yield valid decompositions (Wang et al., 2022; Lim et al., 2022). Such sign and basis ambiguity may cause inconsistency in PEs, leading to unstable representation learning in GNNs. Therefore, models should be invariant to these ambiguities for robust performance.
4. Proposed Framework: GraphPAE
In this section, we introduce GraphPAE, a novel GAE designed to reconstruct both the node features and positions. There are three key components in GraphPAE: corruption, encoder, and decoder, as illustrated in Figure 2(a).
4.1. Data Corruption
The masking-predicting paradigm has become a basic component of autoencoders. A well-designed masking strategy can prevent information leakage and enhance model efficiency (He et al., 2022). Existing GAEs commonly apply the random masking strategy to the discrete nodes and edges. However, node positions, such as eigenvectors, are continuous by nature and are not suitable for the random masking approach. Therefore, we adopt different corruption strategies for these two modalities.
Mask Feature. We follow GraphMAE (Hou et al., 2022) and replace the masked node features with a learnable vector. Specifically, we randomly sample a subset of nodes $\widetilde{\mathcal{V}}\subset\mathcal{V}$ and reset their features. The corrupted feature matrix $\widetilde{\mathbf{X}}$ is defined as:
(1) $\widetilde{\mathbf{X}}_{i}=\begin{cases}\textbf{x}_{[M]},&\text{if }v_{i}\in\widetilde{\mathcal{V}}\\ \mathbf{X}_{i},&\text{if }v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$
where $\mathbf{X}_{i}\in\mathbb{R}^{d}$ is the original feature of node $v_{i}$, and $\textbf{x}_{[M]}\in\mathbb{R}^{d}$ is a learnable vectors.
Offset Position. Inspired by the recent advances in molecular representation learning (Zhou et al., 2023), we propose to add random offsets to node positions. Specifically, we use the top-$k$ eigenvectors of graph Laplacian $\textbf{U}\in\mathbb{R}^{N\times K}$ as the initialization of node positions. Similar to the feature corruption, we define the corruption of node positions as:
(2) $\widetilde{\mathbf{U}}_{i}=\begin{cases}\mathbf{U}_{i}+\delta,&\text{if }v_{i}\in\widetilde{\mathcal{V}}\\ \mathbf{U}_{i},&\text{if }v_{i}\notin\widetilde{\mathcal{V}}\\ \end{cases}$
where $\mathbf{U}_{i}\in\mathbb{R}^{K}$ denotes the position for node $v_{i}$ and $\delta\in\mathbb{R}^{K}$ is a noise vector sampled from a uniform distribution $\mathcal{U}\left(-\mu_{p},\mu_{p}\right)$. In practice, we set $\mu_{p}$ to 0.001 or 0.01 for different datasets.
Relative Positional Encoding. Reconstructing eigenvectors enables GAEs to learn different frequency information. However, it is challenging to recover the corrupted eigenvectors as they suffer from the sign- and basis-ambiguity (Lim et al., 2022; Bo et al., 2024). To address this, we compute the Euclidean distance between each pair of nodes to obtain their relative node distances
(3) ${\mathbf{P}_{i,j}}=\begin{cases}\|\mathbf{U}_{i}-\mathbf{U}_{j}\|_{2},&\text{if }\textbf{A}_{i,j}=1\\ 0,&\text{otherwise}\end{cases}$
where $\mathbf{P}\in\mathbb{R}^{N\times N}$ is the relative distance matrix and we use $\widetilde{\mathbf{P}}$ to indicate its corrupted version. The relative distance matrix is used as a surrogate of node positions to eliminate the ambiguity of eigenvectors.
4.2. GraphPAE Encoder
After corruption, the node features and pair-wise distances are then fed into the encoder to learn node and position representations through message-passing, which can be formulated as
(4) $\mathbf{X}_{i}^{\left(l+1\right)},\mathbf{P}_{i}^{\left(l+1\right)}=f_{\text{enc}}^{\left(l+1\right)}\left(\mathbf{X}_{i}^{\left(l\right)},\left\{\mathbf{X}_{j}^{\left(l\right)}\right\}_{j\in\mathcal{N}_{i}},\mathbf{P}_{i}^{\left(l\right)}\right),$
where $l\in\{0,1,2,...,L\}$ indicates the layer of encoder, and $\mathcal{N}_{i}$ is the neighbors of node $v_{i}$. The first layer of the encoder, e.g., $f_{\text{enc}}^{(0)}$, is designed to align the dimensions of node and position representations. Specifically, it first lifts the scalar relative distance $\mathbf{P}_{i,j}\in\mathbb{R}$ into a vector representation $\mathbf{P}_{i,j}^{(0)}\in\mathbb{R}^{d_{h}}$ through a series of Gaussian RBF kernels
(5) $\mathbf{P}_{i,j}^{(0)}=\text{MLP}\left(\left[G(\mathbf{P}_{i,j};\mu_{1},\sigma),\cdots,G(\mathbf{P}_{i,j};\mu_{d},\sigma)\right]\right),$
where MLP stands for Multi-layer Perceptron and $G(\mathbf{P}_{i,j};\mu_{k},\sigma)=\exp\left(-\left(\mathbf{P}_{i,j}-\mu_{k}\right)^{2}/2\sigma^{2}\right)$ is the $k$-th Gaussian basis functions with mean $\mu_{k}$ and standard deviation $\sigma$. As for the node features, it uses another MLP to transform them into $d_{h}$-dimension representations
(6) $\mathbf{X}_{i}^{\left(0\right)}=\text{MLP}\left(\mathbf{X}_{i}\right).$
After transformation, the encoder needs to aggregate information from both sides to update the node and position representations layer by layer, as shown in Figure 2(b).
Feature Path: PE-enhanced MPNNs. It is well-established that the expressive power of traditional MPNNs is bounded by the 1-WL test (Dwivedi et al., 2023; Cantürk et al., [n. d.]). Fortunately, existing methods prove that adding position information to the message-passing process can significantly improve the expressive power of MPNNs (Dwivedi et al., 2021; Huang et al., 2023). Inspired by the recent progress in graph PEs (Wang et al., [n. d.]), we propose to incorporate the position representation into MPNNs as follows:
(7) $\alpha_{i,j}^{\left(l\right)}=f_{\text{att}}\left(\mathbf{X}_{i}^{\left(l\right)},\mathbf{X}_{j}^{\left(l\right)}\right),\quad\alpha_{i,j}^{\left(l\right)}\in\mathbb{R}^{d},$
$\mathbf{X}_{i}^{\left(l+1\right)}=\sum_{j\in\mathcal{N}_{i}}\left(\alpha_{i,j}^{\left(l\right)}+\mathbf{P}_{i,j}^{\left(l\right)}\right)\odot\text{MLP}\left(\mathbf{X}_{j}^{\left(l\right)}\right),$
where $f_{\text{att}}$ is the attention function to calculate the weights of neighbors and $\odot$ indicates the element-wise multiplication. In general, there are many MPNNs to implement the attention function. For example, if the encoder is GAT (Veličković et al., 2018a), then $d_{h}$ corresponds to the number of attention heads, and the attention function is defined as
(8) $\alpha_{i,j}^{\left(l\right)}=\text{LeakyReLU}\left(\mathbf{w}^{T}\left[\mathbf{W}^{\left(l\right)}\mathbf{X}_{i}^{\left(l\right)}\parallel\mathbf{W}^{\left(l\right)}\mathbf{X}_{j}^{\left(l\right)}\right]\right),\quad\mathbf{w}\in\mathbb{R}^{2d_{h}}.$
For GatedGCN, $d_{h}$ is equal to the dimension of the node representation, and the attention function is defined as
(9) $\alpha_{i,j}^{\left(l\right)}=\text{Sigmoid}\left(\mathbf{W}_{1}^{\left(l\right)}\mathbf{X}_{i}^{\left(l\right)}+\mathbf{W}_{2}^{\left(l\right)}\mathbf{X}_{j}^{\left(l\right)}\right).$
Without loss of generality, we omit edge features in the formulation. Regarding normalization function in Figure 2(b), GAT uses Softmax function, whereas GatedGCN applies degree normalization.
Position Path: Refine Node Positions. The feature path improves the expressiveness of GraphPAE, but still lacks diverse frequency information. To solve this issue, the position path uses the learned attention weights to update the position representations
(10) $\mathbf{P}_{i,j}^{\left(l+1\right)}=\alpha_{i,j}^{\left(l\right)}+\mathbf{P}_{i,j}^{\left(l\right)}.$
Intuitively, the attention weights gradually refine position representations to approach the ground-truth node positions. Notably, the relative distance $\mathbf{P}$ is calculated based on the eigenvectors of graph Laplacian, which contains various frequency information, from low-frequency $\mathbf{u}_{1}$ to high-frequency $\mathbf{u}_{K}$. Therefore, leveraging the refined position representations to reconstruct the relative distance can force the encoder to learn diverse frequency information. We find that this design is quite useful in the heterophilic node classification task, where high-frequency information dominates the classification performance. Experiments can be seen in Section 5.1.
4.3. GraphPAE Decoder
So far, we have described the encoder of GraphPAE, which effectively learns both node and position representations. The decoder then utilizes these representations to recover the corrupted information. In practice, we find that corrupting both node features and positions simultaneously can negatively impact performance, as the recovery process in one path depends on the information from the other. To resolve this, we corrupt only one path’s data during training, leaving the other path’s data intact.
Feature Reconstruction. The node representations are learned by masking node features and preserving the node positions
(11) $\widetilde{\mathbf{X}}_{i}^{(L)},\mathbf{P}_{i}^{(L)}=f_{\text{enc}}\left(\widetilde{\mathbf{X}}_{i}^{\left(0\right)},\left\{\widetilde{\mathbf{X}}_{j}^{\left(0\right)}\right\}_{j\in\mathcal{N}_{i}},\mathbf{P}_{i}^{\left(0\right)}\right),$
To reconstruct the original node features, we apply a feature decoder $f_{\text{dec}}^{X}$ to map the node representations back to the feature space. The reconstruction process is defined as
(12) $\mathbf{X}^{{}^{\prime}}_{i}=f_{\text{dec}}^{X}\left(\widetilde{\mathbf{X}}_{i}^{\left(L\right)}\right),$
where $\mathbf{X}^{{}^{\prime}}_{i}$ is the is the reconstructed feature of node $v_{i}$. We follow GraphMAE and use the scaled cosine error (SCE) as the loss function of feature reconstruction
(13) $\mathcal{L}_{\text{feat}}=\frac{1}{|\widetilde{\mathcal{V}}|}\sum_{v_{i}\in\widetilde{\mathcal{V}}}\left(1-\frac{\mathbf{X}_{i}^{T}\mathbf{X}_{i}^{{}^{\prime}}}{\|\mathbf{X}_{i}\|\cdot\|\mathbf{X}_{i}^{{}^{\prime}}\|}\right)^{\gamma},$
where $\gamma\geq 1$ is a hyperparameter.
Position Reconstruction. The position representations are learned by offsetting eigenvectors and preserving the original node features
(14) $\mathbf{X}_{i}^{(L)},\widetilde{\mathbf{P}}_{i}^{(L)}=f_{\text{enc}}\left(\mathbf{X}_{i}^{\left(0\right)},\left\{\mathbf{X}_{j}^{\left(0\right)}\right\}_{j\in\mathcal{N}_{i}},\widetilde{\mathbf{P}}_{i}^{\left(0\right)}\right),$
Similarly, a position decoder $f_{\text{dec}}^{P}$ is used to recover the original pair-wise node distances
(15) $\mathbf{P}_{i,j}^{{}^{\prime}}=f_{\text{dec}}^{P}\left(\widetilde{\mathbf{P}}_{i,j}^{\left(L\right)}\right)$
For position reconstruction loss, we adopt Huber loss (Huber, [n. d.]), which can make smooth gradients for better convergence
(16) $\mathcal{L}_{\text{pos}}^{i,j}=\begin{cases}\frac{\left(\mathbf{P}_{i,j}^{\prime}-\mathbf{P}_{i,j}\right)^{2}}{2},&\text{if }|\mathbf{P}_{i,j}^{\prime}-\mathbf{P}_{i,j}|<1\\ |\mathbf{P}_{i,j}^{\prime}-\mathbf{P}_{i,j}|-\frac{1}{2},&\text{otherwise}\end{cases}$
$\mathcal{L}_{\text{pos}}=\frac{1}{\sum_{v_{i}\in\widetilde{\mathcal{V}}}\left|\mathcal{N}_{i}\right|}\sum_{v_{i}\in\widetilde{\mathcal{V}},j\in\mathcal{N}_{i}}{\mathcal{L}_{\text{pos}}^{i,j}}$
The overall loss function is formulated as a weighted combination of the feature and position reconstruction losses
(17) $\mathcal{L}=\mathcal{L}_{\text{feat}}+\alpha\mathcal{L}_{\text{pos}}$
where $\alpha$ is the hyperparameter. The pseudocode of GDEM is presented in Algorithm 1.
Algorithm 1 GraphPAE.
1:Input: Graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$, masking ratio $r$, epochs $T$, and noise scale $\mu_{p}$.
2:Preprocess: Compute top-$K$ eigenvectors $\mathbf{U}$ and node distance matrix $\mathbf{P}$ with $\mathbf{U}$.
3:Init: Encoder $f_{\text{enc}}$, feature decoder $f_{\text{dec}}^{X}$, position decoder $f_{\text{dec}}^{P}$, and learnable token $\mathbf{x}_{[M]}$.
4:for $t=1$ to $T$ do
5:     Randomly select $r|\mathcal{V}|$ nodes from $\mathcal{V}$ to form $\widetilde{\mathcal{V}}$.
6:     Replace features in $\widetilde{\mathcal{V}}$ with $\mathbf{x}_{[M]}$ to obtain $\widetilde{\mathbf{X}}$.
7:     Add noise $\delta\sim\mathcal{U}(-\mu_{p},\mu_{p})$ to eigenvectors of nodes in $\widetilde{\mathcal{V}}$ for $\widetilde{\mathbf{U}}$; compute corrupted distances $\widetilde{\mathbf{P}}$ with $\widetilde{\mathbf{U}}$. $\triangleright$ Data Corruption
8:     Encode $\widetilde{\mathbf{X}}$ with $\mathbf{P}$ via $f_{\text{enc}}$ for node representations $\widetilde{\mathbf{X}}^{(L)}$.
9:     Encode $\widetilde{\mathbf{P}}$ with $\mathbf{X}$ via $f_{\text{enc}}$ for positional encodings $\widetilde{\mathbf{P}}^{(L)}$. $\triangleright$ Encoder
10:     Decode $\widetilde{\mathbf{X}}^{(L)}$ and $\widetilde{\mathbf{P}}^{(L)}$ via $f_{\text{dec}}^{X}$ and $f_{\text{dec}}^{P}$ to reconstruct features $\mathbf{X}^{\prime}$ and distances $\mathbf{P}^{\prime}$.
11:     Compute $\mathcal{L}_{\text{feat}}$ with $\mathbf{X}$ and $\mathbf{X}^{\prime}$, and $\mathcal{L}_{\text{pos}}$ with $\mathbf{P}$ and $\mathbf{P}^{\prime}$.
12:     $\mathcal{L}=\mathcal{L}_{\text{feat}}+\alpha\mathcal{L}_{\text{pos}}$. $\triangleright$ Decoder
13:     Update $f_{\text{enc}},f_{\text{dec}}^{X}$, and $f_{\text{dec}}^{P}$ by minimizing $\mathcal{L}$.
14:end for
15:return Trained encoder $f_{\text{enc}}$.
5. Experiments
In this section, we conduct extensive experiments, including unsupervised node classification, unsupervised graph prediction, and transfer learning on large-scale molecule graphs to verify the effectiveness of GraphPAE. Moreover, we perform ablation studies on position reconstruction and dual-path design. Finally, we analyze the influence of the position reconstruction loss.
Table 2. Node classification results of different graph self-supervised learning, mean accuracy (%) ± standard deviation. Bold indicates the best performance and underline means the runner-up.
Dataset Small Graphs Large Graphs
BlogCatalog Chameleon Squirrel Actor arXiv-year Penn94
Supervised 80.52±2.10 80.02±0.87 71.91±1.03 33.93±2.47 46.02±0.26 81.53±0.55
DGI 72.07±0.16 43.83±0.14 34.56±0.10 27.98±0.09 - -
BGRL 79.74±0.46 61.24±1.07 43.24±0.52 26.61±0.57 41.43±0.04 63.31±0.49
MVGRL 63.24±0.94 73.19±0.42 60.09±0.44 34.64±0.20 - -
CCA-SSG 74.00±0.28 75.00±0.75 61.58±1.98 27.79±0.58 40.78±0.01 62.63±0.20
Sp2GCL 72.73±0.46 78.88±1.04 62.61±0.87 34.70±0.92 39.09±0.02 68.80±0.45
VGAE 60.47±1.84 62.32±1.90 42.50±1.35 31.57±0.75 36.39±0.21 55.31±0.28
GraphMAE 79.90±1.13 79.50±0.57 61.13±0.60 32.15±1.33 40.30±0.04 67.97±0.21
GraphMAE2 77.34±0.12 79.13±0.19 70.27±0.88 34.48±0.26 38.97±0.03 67.86±0.42
MaskGAE 73.10±0.08 74.50±0.87 68.53±0.44 33.44±0.34 40.59±0.04 63.84±0.03
S2GAE 75.76±0.43 60.24±0.37 68.60±0.56 26.17±0.38 40.32±0.12 70.24±0.09
AUG-MAE 82.03±0.69 70.10±1.88 62.57±0.67 33.42±0.38 37.10±0.13 69.90±0.43
GraphPAE 85.76±1.22 80.51±1.25 72.05±1.40 38.55±1.35 41.85±0.04 71.79±0.37
Table 3. Graph regression and classification results of different graph self-supervised learning on OGB datasets. Bold indicates the best performance and underline means the runner-up. $\downarrow$ means lower the better and $\uparrow$ means higher the better.
[…]
Dataset molesol molipo molfreesolv molbace molbbbp molclintox moltocx21
Supervised 1.173±0.057 0.757±0.018 2.755±0.349 80.42±0.96 68.17±1.48 88.14±2.51 74.91±0.51
InfoGraph 1.344±0.178 1.005±0.023 10.005±8.147 73.64±3.64 66.33±2.79 64.50±5.32 69.74±0.57
GraphCL 1.272±0.089 0.910±0.016 7.679±2.748 73.32±2.70 68.22±2.19 74.92±4.42 72.40±1.07
MVGRL 1.433±0.145 0.962±0.036 9.024±1.982 74.88±1.43 67.24±3.19 73.84±2.75 70.48±0.83
JOAO 1.285±0.121 0.865±0.032 5.131±0.782 74.43±1.94 67.62±1.29 71.28±4.12 71.38±0.92
Sp2GCL 1.235±0.119 0.835±0.026 4.144±0.573 78.76±1.43 68.72±1.53 80.88±3.86 73.06±0.75
GraphMAE 1.050±0.034 0.850±0.022 2.740±0.233 79.14±1.31 66.55±1.78 80.56±5.55 73.84±0.58
GraphMAE2 1.225±0.081 0.885±0.019 2.913±0.293 80.74±1.53 67.67±1.44 75.75±3.65 72.93±0.69
StructMAE 1.499±0.043 1.089±0.002 2.568±0.262 77.75±0.42 65.66±1.16 79.42±4.56 71.13±0.61
AUG-MAE 1.248±0.026 0.917±0.013 2.395±0.158 78.54±2.49 67.05±0.63 82.66±1.98 74.33±0.07
GraphPAE 1.015±0.045 0.810±0.018 2.058±0.188 81.11±1.24 68.56±0.71 82.69±3.39 74.46±0.54
5.1. Node Classification
Dataset. We evaluate the performance of GraphPAE on 6 representative heterophilic graphs: BlogCatalog (Meng et al., 2019), Chameleon, Squirrel, Actor (Pei et al., 2020), arXiv-year (Hu et al., 2020b), and Penn94 (Traud et al., 2012). Specifically, arXiv-year and Penn94 are large-scale graphs (¿ 40,000) to evaluate the scalability of the methods. For all the datasets, we use the public splits.
Baselines and Settings. We benchmark GraphPAE against a wide range of graph self-supervised baselines, which can be roughly divided into: (1) contrastive learning, i.e., DGI (Veličković et al., 2018b), BGRL (Thakoor et al., 2021), MVGRL (Hassani and Khasahmadi, 2020), CCA-SSG (Zhang et al., 2021b), and $\text{Sp}^{2}\text{GCL}$ (Bo et al., 2024). (2) graph autoencoders, i.e., VGAE (Kipf and Welling, 2016), GraphMAE (Hou et al., 2022), GraphMAE2 (Hou et al., 2023), MaskGAE (Li et al., 2023a), S2GAE (Tan et al., 2023), and AUG-MAE (Wang et al., 2024). We use GAT with 4 heads and 1024 hidden units as the encoder for all methods, and the number of layers is searched in the range of {2, 3}. Moreover, we adopt two-layer MLPs with ReLU activation as both feature and position decoders for efficiency and scalability. In the evaluation protocol, we freeze the encoder and generate node representations. The node representations are then input into a linear classifier for training with labeled data and inference for node classification. For all methods, we use the Adam optimizer and run 10 times on each graph.
Results. As shown in Table 2, GraphPAE consistently outperforms all baseline methods across all datasets, demonstrating the effectiveness of our framework. We highlight two important observations as follows. (1) GraphPAE surpasses both spatial- and spectral-based contrastive learning methods on most datasets, indicating its ability to effectively encode both spatial and spectral patterns through the joint reconstruction of features and positions. (2) GraphPAE also consistently outperforms existing GAEs, including those focused on feature reconstruction, structure reconstruction, and hybrid strategies. This superior performance is mainly attributed to two key factors. First, the incorporation of positional encodings enhances the expressivity of node representations. Second, reconstructing positions encourages the GAE to capture diverse frequency information, leading to high-quality graph representations.
Table 4. Quantum chemistry property results of transfer learning on QM9. The best and runner-up results are highlighted with bold and underline, respectively.
Target $\mu$ $\alpha$ $\epsilon_{\text{homo}}$ $\epsilon_{\text{lumo}}$ $\Delta_{\epsilon}$ $R^{2}$ ZPVE $U_{0}$ $U$ $H$ $G$ $C_{v}$
Unit D $a_{0}^{3}$ $10^{-2}$meV $10^{-2}$meV $10^{-2}$meV $a_{0}^{2}$ $10^{-2}$meV meV meV meV meV cal/mol/K
GraphCL 1.035 2.321 2.030 3.667 4.523 40.725 2.063 2.461 1.745 1.734 1.751 1.747
GraphMAE 1.030 2.924 2.407 6.373 4.813 41.955 4.623 1.411 2.207 2.208 2.207 2.200
Mole-BERT 1.031 1.918 1.477 4.127 4.240 44.374 2.190 2.532 2.509 2.511 2.516 2.508
SimSGT 1.064 2.413 2.837 4.227 4.107 40.504 2.127 1.948 2.420 2.416 2.416 2.410
GraphPAE 0.703 0.879 1.199 2.141 2.289 36.480 0.502 0.510 0.639 0.639 0.641 0.643
Table 5. Ablation studies of position reconstruction and framework design on node- and graph-level tasks. Exp No.: the number of different experimental settings. Corrupt Info.: corrupted information during training. Recon Info.: information to be reconstructed during training. Bold indicates the best performance.
Exp Corrupt Info. Recon Info. Dual-Path Node-level Graph-level
No. Feature Position Feature Position Blog ($\uparrow$) Squirrel ($\uparrow$) Bace ($\uparrow$) Bbbp ($\uparrow$) Freesolv ($\downarrow$)
a 82.8±1.7 66.4±1.6 78.4±1.2 66.4±1.7 2.79±0.40
b 83.5±1.0 68.5±0.9 78.9±2.1 66.8±0.6 2.44±0.36
c 84.6±1.6 71.3±0.9 79.4±3.4 67.7±0.9 2.20±0.14
d 85.8±1.2 72.1±1.4 81.1±1.2 68.6±0.7 2.06±0.19
5.2. Graph Prediction
Datasets. For graph-level tasks, we evaluate GraphPAE on 7 OGB datasets (Hu et al., 2020b), including 3 graph regression tasks and 4 graph classification tasks. For all datasets, we use public splits for a fair comparison. We use MSE and ROC-AUC as the evaluation metrics for regression and classification, respectively.
Baselines and Settings. We select 5 contrastive learning methods, i.e., InfoGraph (Sun et al., 2019), GraphCL (You et al., 2020), MVGRL (Hassani and Khasahmadi, 2020), JOAO (You et al., 2021), and $\text{Sp}^{2}\text{GCL}$ (Bo et al., 2024), and 4 graph autocoders, i.e., GraphMAE (Hou et al., 2022), GraphMAE2 (Hou et al., 2023), StructMAE (Liu et al., 2024b), and AUG-MAE (Wang et al., 2024). We use a two-layer GatedGCN as the encoder and set the hidden dimension $d_{h}=300$. We adopt two-layer MLPs as feature and position decoders. For evaluation, we freeze the encoder to output node representations and input them into pooling functions for graph representations. Similarly, we input the graph representations into a linear classifier to evaluate the performance for downstream tasks. We use Adam optimizer and report the metrics with mean results and standard deviation of 10 seeds.
Results. Table 3 summarizes the results on graph-level tasks. GraphPAE consistently achieves superior performance across both regression and classification benchmarks, highlighting its ability to learn high-quality graph representations. Notably, GraphPAE shows competitive performance against contrastive learning methods and even exhibits notable improvements compared against the spectral-based method $\text{Sp}^{2}\text{GCL}$ on molfreesolv, molbace, and molclintox. In addition, GraphPAE also outperforms other GAEs across most datasets. We attribute these gains to the proposed position corruption-reconstruction strategy, which improves the encoder’s ability to identify crucial substructures for downstream tasks.
5.3. Transfer Learning
Settings. We conduct transfer learning experiments on molecule property prediction to evaluate the generalization ability of GraphPAE. Specifically, we follow the setting of (Liu et al., 2024a), which first pre-trains the encoder on 2 million molecules sampled from ZINC15 (Sterling and Irwin, 2015), and then fine-tunes on QM9 (Wu et al., 2018) to predict the quantum chemistry properties. We compare GraphPAE with the state-of-the-art molecule graph pre-training model, i.e., GraphCL (You et al., 2020), GraphMAE (Hou et al., 2022), Mole-BERT (Xia et al., 2023), and SimSGT (Liu et al., 2024a). We use a five-layer GatedGCN with 300 hidden units. Upon finishing pre-training, a two-layer MLP is attached after the graph representations for property prediction. During the fine-tuning protocol, the encoder and attached MLP are trained together on labeled data of downstream tasks. Consistent with (Liu et al., 2024a), we divided QM9 into train/validation/test sets with 80%/10%/10% by scaffold splits. We use Adam optimizer and run 5 times to report average MAE and standard deviation. The results are reported in Table 4.
Results. We observe that GraphPAE demonstrates robust generalization across all prediction targets compared to state-of-the-art models. Among the baselines, Mole-BERT and SimSGT are specially designed for molecular graph pre-training, incorporating customized reconstruction objectives tailored to the unique characteristics of molecular graphs. Specifically, Mole-BERT pretrains a discrete codebook within the subgraph of nodes, inherently capturing both node features and local structural patterns. During reconstruction, the model replaces the raw features with codebook entries as predictive targets. SimSGT employs a tokenizer to encode feature and local structure information, using the tokenized outputs as reconstruction targets. Despite these tailored strategies, GraphPAE consistently outperforms these models. The superiority is largely attributed to the effective positional encodings obtained by position corruption-reconstruction, enabling the graph representations to capture crucial substructures for prediction targets such as cycles.
5.4. Ablation Studies
To verify the effectiveness of important designs of GraphPAE, we conduct extensive ablation experiments. We select 2 datasets for node-level tasks and 3 datasets for graph-level tasks.
Effectiveness of Position Reconstruction. To achieve position corruption-reconstruction, GraphPAE inevitably introduces positional encodings into the encoder. Firstly, we verify that the superior performance is attributed not only to the introduction of positional encoding but also to the position reconstruction. We conducted two sets of comparative experiments in Table 5: one comparing Exp a with Exp b, and another comparing Exp c with Exp d. We have the following summaries: (1) In Exp a, both features and positions are corrupted, but only the features are reconstructed. In Exp b, both features and positions are reconstructed. We observe that Exp b performs better than Exp a, demonstrating the effectiveness of position reconstruction. (2) In Exp c, we remove both position corruption and reconstruction from GraphPAE. Comparing Exp d against Exp c, we find that the encoder incorporating position reconstruction consistently improves performance over that solely with feature reconstruction.
Effectiveness of Dual-Path Reconstruction. Comparing Exp b and Exp c in Table 5, we observe that position reconstruction without the dual-path architecture fails to improve performance and may even lead to degradation. Specifically, in Exp b, both masked features and noisy positions are reconstructed directly from node representations. However, recovering relative distances from node representations generated with corrupted positions is more difficult than recovering them from refined distance encoding directly. Rather than enhancing node representations or capturing diverse frequency information, the corrupted positional inputs introduce additional noise during training. As a result, applying position corruption-reconstruction without appropriate design hinders the model’s ability to learn meaningful structural patterns. Based on these observations, we propose a dual-path architecture that disentangles feature and relative distance encodings for reconstruction. Experimental results from Exp d demonstrate that this strategy consistently achieves the best performance across datasets, validating the effectiveness of our dual-path encoder design.
Improvement Attributed to $\mathcal{L}_{pos}$ and PEs. Since integrating positional encodings also enhances representation learning, it remains unclear how much of the performance gain in GraphPAE stems from the position corruption-reconstruction mechanism, and how much from the positional information itself. Therefore, we conduct ablation studies on both $\mathcal{L}_{pos}$ and positional encodings. Specifically, we remove $\mathcal{L}_{pos}$ from Equation 17 while keeping PEs, to evaluate the effect of position reconstruction. Then, we remove both $\mathcal{L}_{pos}$ and PEs to isolate the impact of positional information. The results are reported in Table 6. Notably, even with positional encodings retained, removing $\mathcal{L}_{pos}$ consistently leads to a noticeable performance decline across all datasets, highlighting the importance of the proposed position corruption-reconstruction mechanism in enhancing GAEs.
Table 6. Ablation studies of $\mathcal{L}_{pos}$ and positional encodings.
Methods Bace ($\uparrow$) Bbbp ($\uparrow$) Freesolv ($\downarrow$)
GraphPAE 81.11±1.24 68.56±0.71 2.058±0.19
w/o $\mathcal{L}_{pos}$ 79.40±3.45 67.74±0.92 2.204±0.14
w/o $\mathcal{L}_{pos}$ PEs 79.14±1.31 66.55±1.78 2.740±0.23
(a) Molbbbp
(b) BlogCatalog
(c) Squirrel
(d) Chameleon
Figure 3. Influence of the loss weight $\alpha$ and the number of eigenvectors $K$.
5.5. Parameters Analysis
We conduct additional parameter analysis of the loss weight $\alpha$ and the number of eigenvectors $K$.
Influence of the loss weight $\alpha$ in GraphPAE. Figure 3(a) and 3(b) presents the hyperparameter analysis of $\alpha$ to further examine the influence of the positional reconstruction loss $\mathcal{L}_{pos}$. We summarize the key observations as follows: (1) As $\alpha$ increases within a certain threshold, the performance improves progressively, indicating that position reconstruction contributes positively to the quality of learned representations. (2) However, excessively large values of $\alpha$ lead to performance degeneration. This suggests assigning a high weight to position reconstruction may cause the encoder to overemphasize positional information at the expense of overall representation learning. (3) While the optimal value of $\alpha$ varies slightly across datasets, we observe that the best-performing range typically lies within [1e-3, 1e-2, 1e-1], making it easy to search in practice. Additional hyperparameter details are provided in Appendix A.4.
Influence of the number of eigenvectors $K$ in GraphPAE. We further analyze the effect of the number of eigenvectors $K$ on performance using the Squirrel and Chameleon datasets, as shown in Figure 3(c) and 3(d). As $K$ increases, performance generally improves due to the model’s ability to capture a wider range of frequency information. However, once $K$ surpasses a certain threshold, the performance gains become marginal. Therefore, to balance effectiveness and computational efficiency, we adopt a moderate value of $K$, typically in the range of [50-100], which provides a good trade-off in practice.
6. Conclusion
In this paper, we propose GraphPAE, a graph autoencoder that reconstructs both node features and relative distance. To enhance models’ expressivity to distinguish intricate patterns and the ability to integrate various frequency information, GraphPAE introduces position corruption and recovery into GAEs and designs a dual-path reconstruction strategy. Extensive experiments, including node classification, graph prediction, and transfer learning, demonstrate the superiority of our GraphPAE.
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China (No. 62192784, U22B2038, 62472329).
Appendix A Experimental Details
A.1. Spectral Analysis of Masking Strategies
Masking Strategies for Node Features and Edges. We conduct the experiments in the Squirrel datasets with 20% masking ratio. For node feature masking, given the node feature matrix $\mathbf{X}$, we randomly mask 20% of the node in the dataset and set the feature vector to zero. The remaining 80% of nodes retain their original feature. The corrupted feature matrix is defined as $\widetilde{\mathbf{X}}$. For edge masking, given the original graph structure $\mathbf{A}$, we randomly remove 20% of the edges from the graph while preserving the node features. This process results in a structurally corrupted graph with a new adjacency matrix $\widetilde{\mathbf{A}}$.
Computation of Frequency Magnitudes. First, we calculate Laplacian matrix by:
(18) $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
The eigenvectors of $\mathbf{L}$ are decomposed as $\mathbf{L}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top}$, where $\bm{\Lambda}=\text{diag}(\{\lambda_{i}\}_{i=1}^{N})$, $\mathbf{U}=[\mathbf{u}_{1},\cdots,\mathbf{u}_{N}]\in\mathbb{R}^{N\times N}$, and the eigenvalues are ordered as $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$. $\lambda_{i}$ reflects the frequency magnitude of corresponding eigenvector $\mathbf{u}_{i}$ over the graph. We transform node features into the spectral domain by
(19) $\mathbf{X}^{s}=\mathbf{U}^{\top}\mathbf{X},\quad\mathbf{X}^{s}\in\mathbb{R}^{N\times d}.$
In particular, the i-th row of $\mathbf{X}^{s}$ corresponds to the frequency magnitude at eigenvalue $\lambda_{i}$. With $d$-dimensional $\mathbf{X}^{s}$, we first calculate the mean magnitude by:
(20) $\bar{\mathbf{X}}^{s}=\frac{1}{d}\sum_{i=1}^{d}\mathbf{X}_{:,i}^{s},\quad\bar{\mathbf{X}}^{s}\in\mathbb{R}^{N}$
For a given frequency band $[f_{1},f_{2}]$, we extract the set of indices $\mathcal{I}$ corresponding to eigenvalues within this range:
(21) =$\mathcal{I}=\{i\mid f_{1}\leq\lambda_{i}\leq f_{2}\}$
The average frequency magnitude within this band is computed as:
(22) $\bar{\mathbf{X}}^{s}_{[f_{1},f_{2}]}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\bar{\mathbf{X}}^{s}[i],\quad\bar{\mathbf{X}}^{s}_{[f_{1},f_{2}]}\in\mathbb{R}$
For the feature masking case, we calculate $\mathbf{L}$ by substituting $\mathbf{A}$ into Equation 18. Then we decompose $\mathbf{L}$ and get $\mathbf{U}$. By substituting $\mathbf{U}$, $\mathbf{X}$ and $\mathbf{U}$, $\widetilde{\mathbf{X}}$ in Equation 19-22, we can get the frequency magnitude of original and corrupted graphs. The comparison of differences is shown in Figure 1(a).
For edge masking case, we calculate $\mathbf{L}$ and $\widetilde{\mathbf{L}}$ by substituting $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ into Equation 18. Then we decompose $\mathbf{L}$ and $\widetilde{\mathbf{L}}$, and get $\mathbf{U}$ and $\widetilde{\mathbf{U}}$. By substituting $\mathbf{U}$, $\widetilde{\mathbf{U}}$ and $\mathbf{X}$ in Equation (19) - (22), we can also get the frequency magnitude of original and corrupted graphs as shown in Figure 1(b).
A.2. Additional Experiment Results.
We further evaluate GraphPAE on homophilic graphs Facebook (Rozemberczki et al., 2021) and Wiki (Mernyei and Cangea, 2020), as well as a large-scale graph ogbn-Products (Hu et al., 2020b). Notably, ogbn-Products contains 2,449,029 nodes and 61,859,140 edges, making it a suitable benchmark to demonstrate the scalability of GraphPAE. The experimental results are summarized in Table 7, and we make the following observations: (1) GraphPAE consistently achieves competitive performance across these different benchmarks. (2) The performance gains on homophilic graphs are less pronounced compared to those on heterophilic graphs. We conjecture this is because, in homophilic graphs, a small portion of very low-frequency information is sufficient.
Table 7. Additional experiment results on homophilic and large-scale graphs.
Methods Facebook Wiki ogbg-Products
BGRL 89.71±0.35 79.02±0.13 78.59±0.02
CCA-SSG 89.45±0.60 78.85±0.32 75.27±0.05
GraphMAE 89.54±0.36 78.94±0.48 78.89±0.01
GraphMAE2 88.49±0.43 78.84±0.44 81.59±0.02
GraphPAE 91.46±0.23 79.32±0.29 79.10±0.02
A.3. Statistics of Datasets
We benchmark GraphPAE across various tasks, including node classification, graph prediction, and transfer learning. Specifically, we conduct experiments on six node classification datasets and seven graph-level prediction datasets and use ZINC-2M and QM9 for transfer learning. For datasets in node classification and graph prediction, we adopt public data splits. In transfer learning, we follow (Liu et al., 2024a) and divide QM9 into train/valid/test sets with 80%/10%/10% by scaffold split. Detailed statistics of the node- and graph-level datasets are presented in Table 8 and Table 9, respectively.
A.4. Hyperparamters
Hyperparameter details of node classification and graph prediction are reported in Table 10 and Table 11, respectively. $r$ denotes the mask ratio and $\alpha$ is the loss weight of $\mathcal{L}_{pos}$. dp and $\text{dp}_{edge}$ represent the dropout of node features and edge features, respectively. For transfer learning, we pre-train GraphPAE on 2 million molecules sampled from ZINC15 with 100 epochs. The mask ratio is set to 0.35, and the loss weight for $\mathcal{L}_{pos}$ is 0.01. Both node and edge dropout are set to 0.0, and the number of eigenvectors for positional encoding is 6. During fine-tuning, the encoder and the attached MLP are jointly trained on the QM9 dataset using an initial learning rate of 0.001 and no weight decay. We apply a dropout rate of 0.1 to both node features and edge features within GNN layers.
A.5. Complexity Analysis
The eigen-decomposition complexity is $\mathcal{O}(N^{3})$. However, we only need to decompose the smallest $K$ eigenvalues, which reduces complexity to $\mathcal{O}\left(N^{2}K\right)$. Additionally, the decomposition is performed once per graph, so its cost is amortized over the entire experiment.
During training, GraphPAE introduces an additional $\mathcal{O\left(\mathcal{E}\right)}$ complexity per layer for PE computation and an extra cost for relative node distance reconstruction. Since we only reconstruct corrupted node pairs, the complexity remains much lower than $\mathcal{O}\left(\mathcal{E}\right)$. During inference, the extra overhead also comes from PE computation.
Overall, GraphPAE introduces an additional $\mathcal{O}(\mathcal{E})$ complexity in training. However, this slight overhead enables the encoder to capture more diverse frequency information.
Table 8. Statistics of node classification datasets.
Datasets Nodes Edges Features Classes Train/Valid/Test
BlogCatalog 5,196 343,486 8,189 6 120/1,000/1,000
Chameleon 2,277 62,792 2,325 5 1,092/729/456
Squirrel 5,201 396,846 2,089 5 2,496/1,664/1,041
Actor 7,600 53,411 932 5 3,648/2,432/1,520
arXiv-year 169,343 2,315,598 128 5 84671/42335/42337
Penn94 41,554 2,724,458 4,814 2 19,407/9,703/9,705
Table 9. Statistics of graph prediction datasets.
Dataset Graphs Avg. Nodes Avg. Edges Classes Task Metric
molesol 1,128 13.3 13.7 1 Regression RMSE
mollipo 4,200 27.0 29.5 1 Regression RMSE
molfreesolv 642 8.7 8.4 1 Regression RMSE
molbace 1,513 34.1 36.9 1 Binary Class. ROC-AUC
molbbbp 2,039 24.1 26.0 1 Binary Class. ROC-AUC
molcintox 1,477 26.2 27.9 2 Binary Class. ROC-AUC
moltox21 7,831 18.6 19.3 12 Binary Class. ROC-AUC
ZINC-2M 2,000,000 26.62 57.72 - Pre-Training -
QM9 1,177,631 8.80 18.81 12 Targets Finetuning MAE
Table 10. Hyperparameters of node classification.
Dataset lr wd $r$ $\alpha$ dp $\text{dp}_{\text{edge}}$ $K$
BlogCatalog 0.001 0.0 0.25 0.1 0.6 0.0 50
Chameleon 0.001 0.0 0.25 0.01 0.0 0.0 50
Squirrel 0.001 0.0 0.5 0.001 0.6 0.0 50
Actor 0.0005 0.0 0.25 0.01 0.0 0.0 50
arXiv-year 0.001 0.0 0.5 0.01 0.0 0.0 100
Penn94 0.001 0.0 0.25 0.001 0.0 0.0 200
Table 11. Hyperparameters of graph prediction.
Dataset pooling epoch lr wd $r$ $\alpha$ dp $\text{dp}_{\text{edge}}$ $K$
molesol sum 20 0.0005 0.0 0.75 0.1 0.6 0.5 8
molipo sum 20 0.0005 0.0001 0.25 0.001 0.6 0.0 30
molfreesolv sum 100 0.0001 0.0 0.5 0.1 0.5 0.5 15
molbace mean 100 0.001 0.0 0.75 0.1 0.5 0.5 30
molbbbp mean 20 0.001 0.0 0.5 0.01 0.6 0.6 30
molclintox mean 20 0.0001 0.0 0.25 0.01 0.6 0.0 30
moltocx21 mean 20 0.0001 0.0 0.25 0.1 0.0 0.6 8